OpenStudy (ny,ny):

Find the vertical asymptotes, holes, intercepts, end behavior, and graph for the function y= x(x-1) / x^2-1.

OpenStudy (mathmale):

There'd be less ambiguity here if you'd please enclose the denominator "x^2-1" inside parentheses. Want to emphasize that the numerator, x(x-1), is divided by all of "x^2-1." You will be asked repeatedly to find asymptotes, intercepts, end behavior, etc., so it'd be worth the investment of your time and effort to learn these now. Hint #1: a VERTICAL INTERCEPT is found by letting x=0. If the function is defined at x=0, the numeric value that results is the y-coordinate of the y-intercept (or vertical intercept).

OpenStudy (mathmale):

What specifically do you need to know to complete this problem?

OpenStudy (ny,ny):

Ok, so to find the y intercept: y= x(x-1) / (x^2-1) y= (0) / (-1) y=0 How do I find the x intercept? Is it zero because thats the only value to make it zero? How do I know where is the vertical asymptote? I know the domain is x cannot equal 1, and -1 I think the hole is at x=1 because it can cancel on the top and bottom Um end behavior is y=1? And how to graph based on this?

OpenStudy (triciaal):